Circular No. 9048

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2009 K4 (GIBBS) A. R. Gibbs reports his discovery of a comet on unfiltered CCD images

obtained at low altitude in astronomical twilight with the Catalina 0.68-m Schmidt telescope (discovery observation tabulated below); four co-added 15-s exposures show a 5′′ × 9′′ coma, elongated in the direction of motion, with a “fuzzy halo” 11′′ across. Four co-added 30-s follow-up images obtained by R. E. Hill on May 28.2 UT with the 1.5-m reflector at Mt. Lemmon show a coma 20′′–24′′ in diameter with no obvious tail. After posting on the Minor Planet Center’s ‘NEOCP’ webpage, other CCD astrometrists have also commented on the object’s cometary appearance. J. Young (Table Mountain 0.61-m reflector, May 27.2 and 28.2, low altitude; nearby faint galaxies were problematical) reports a very diffuse coma of diameter 12′′ with little central condensation and no tail. M. Pietschnig (Alter Satzberg, Vienna, Austria, 0.35-m reflector, May 27.84–27.86) measures a 15′′ coma. M. Jäger (Stixendorf, Austria, 0.20-m reflector, May 27.9) reports a coma of diameter 1′–1′.5 of total mag ∼ 16 with a nuclear condensation of mag ∼ 17.5. E. Guido and G. Sostero, Castellammare di Stabia, Italy, report that twenty-one co-added 120-s unfiltered images obtained remotely on May 28.2 with a 0.25-m reflector near Mayhill, NM, U.S.A., show a sharp central condensation ∼ 6′′ in diameter surrounded by an extremely faint outer coma nearly 40′′ in diameter. C. Hergenrother, Lunar and Planetary Laboratory, reports that a co-added 11-min R-band exposure with the Kuiper 1.54-m reflector shows a circular coma 17′′ in diameter.

2009 UT α_{2000} δ_{2000} Mag. Observer
May 27.14873 7^{h}57^{m}43.^{s}63 +46^{o}02′11′′.6 17.0 Gibbs

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2009-K63.

T = 2009 June 18.433 TT  ω = 126.^{o}837
Ω = 29.747 }2000.0
q = 1.56252 AU  i = 34.820

V5581 SAGITTARII = NOVA SAGITTARII 2009 No. 1 N. Samus reports that the GCVS team has assigned the designation

V5581 Sgr to the nova announced on IAUC 9041.

2009 May 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green